OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3

OpenStudy (aroub):

You find the slope using the slope formula which is: y2-y2/x2-x1 = 8-2/-5+3 = 6/-2 = -3 And now you have to find the y-intercept: y=mx+b; (-3,2) , m=-3 2=-3(-3)+b 2=9+b -7=b => y=-3x-7